# Instructions for EMNLP 2023 Proceedings

### v1

@@ Abstract
When engaging in collaborative tasks, humans efficiently exploit the semantic structure of a conversation to optimize verbal and nonverbal interactions. But in recent “language to code" or “language to action" models, this information is lacking. We show how incorporating the prior discourse and nonlinguistic context of a conversation situated in a nonlinguistic environment can improve the “language to action" component of such interactions. We fine tune an LLM to predict actions based on prior context; our model, NeBuLa, doubles the net-action F1 score over the baseline on this task of Jayannavar et al., (2020). We also investigate our model’s ability to construct shapes and understand location descriptions using a synthetic dataset.
@@ 1 Introduction
The first factor is the interactions between linguistic and nonlinguistic contexts. Previous work has shown that at least some context is needed to understand and carry out conversationally given instructions Jayannavar et al., (2020). We improve on that work by first establishing a baseline by using the entire exchange up to an instruction $i$ as a context for an LLM to interpret $i$. Our LLM model, NeBuLa (Neural Builder with Llama), trained on the Minecraft Dialogue Corpus (MDC) Narayan-Chen et al., (2019), achieves net-action F1 scores that is almost double of Jayannavar et al., (2020). Using the Minecraft Structured Dialogue dataset (MSDC) Thompson et al., (2024), which provides semantic relations between MDC dialogue moves and nonlinguistic actions, we show that particular discursive components of the linguistic and nonlinguistic context are necessary and sufficient for the LLM to understand an instruction to the degree provided by the baseline.
@@
Analysing NeBuLa’s performance revealed two other factors that importantly adversely affect its performance. An instruction in the MSDC has two basic components: a description of a shape in terms of four parameters—numbers of components, colors, arrangement and orientation— and the description of a location where the shape should be placed. Human Architects often use analogies to everyday objects that may be challenging to process, and in addition shape descriptions are often underspecified, meaning that one could perform the instruction correctly in various ways. Location descriptions in the Minecraft world are also quite difficult to process and highly underspecified. For example, put a tower in a corner could be correctly located in any of the four corners of the Minecraft board. We address this problem in two ways: first by further finetuning NeBuLa on a synthetic dataset to improve its performance in building basic shapes and locating them appropriately; and secondly, and more importantly, by revising the evaluation metric used by Jayannavar et al., (2020) to reflect more realistically the semantics of location expressions. We show that, on our synthetic dataset, NeBuLa achieves high accuracy as per our intuitive metric in performing basic instructions.
@@
After some preliminaries and discussion of prior work (Section 2), we present the NeBuLa model and its baseline performance in Section 3, and then a necessary and sufficient discourse feature to get scores equivalent to the baseline in Section 4. In Section 5, we explain several issues associated with Minecraft corpus. We try to address these issues in Section 6. In this section, we explain our evaluation metric for underspecified instructions, as well as experiments on our synthetic datasets.
@@ 2 Related Work
MDC Narayan-Chen et al., (2019) introduced a corpus of two person dialogues situated in a simulated Minecraft environment. The dialogues record conversations about collaborative tasks, in which an Architect and a Builder cooperate to build sometimes complex 3-dimensional shapes out of blocks of six different colors. The Architect provides instructions, while the Builder is tasked with translating these instructions into actions. The Builder sometimes asks questions, and the Architect may correct themselves or the Builder, or both, concerning both linguistic and nonlinguistic moves. The corpus accurately reflects the variety and complexity of actual cooperative conversation. Details on the MDC are in Table 1.
@@
Instructions to code: Neural Builder and variants The MDC Jayannavar et al., (2020) incentivized the development of an algorithm that could predict sequences of actions from instructions. The actions involved basic moves of placing or removing blocks from certain positions in the environment. Jayannavar et al., (2020) trained a model consisting of a GRU Cho et al., (2014) to handle textual input coupled with a CNN to integrate information from the current state and a GRU to predicted an action sequence. Although they experimented with several training regimes, the best performance came from one in which a sequence of conversational moves after some action sequence, assumed to be instructions are given to the model, are followed by the next action sequence of the Builder, followed by the next sequence of linguistic moves are input to the model to predict the subsequent action sequence. (See Figure 1).
@@
Figure 1: The Neural Builder Jayannavar et al., (2020) takes as input the sequence $i\_n$ $a\_n$ $i\_n+1$ and the world-state to predict the subsequent action sequence.
@@
The net-action F1 metric evaluates a model’s prediction based on the exact color and coordinate match between the model’s predicted sequence and Builder’s gold action sequence. In general, Jayannavar et al., (2020) showed that the problem of predicting action sequences from natural language instructions in naturally occurring dialogue remains extremely challenging. Their Neural Builder had net action f1 of $0.20$ on the MDC test set.
@@
Shi et al., (2022) propose a somewhat different task from Jayannavar et al., (2020); they try to predict when the Builder should execute an action and when they should instead ask for a clarification question. To this end, they annotated all Builder dialogue moves with a taxonomy of dialogue acts. They then specified a single specific action under the execution label instead of a sequence of actions. Thus, their set-up is not directly comparable to that of Jayannavar et al., (2020).
@@
Bonial et al., (2020, 2021) added dialogue acts to Minecraft utterances, but they did not evaluate the effect of these dialogue acts on the Neural Builder’s predictions of actions. Dialogue acts are a partial step towards a full discourse structure; they provide labels for various dialogue moves, but the full discourse structure that we propose to use involves relations between moves. These relations are important as they tell us how to link different parts of, for instance, an instruction into a coherent whole. As we aim to demonstrate in this paper, discourse structure can help to clean up datasets for training and thereby improve training.
@@
MSDC Thompson et al., (2024) provided full discourse annotations for the Minecraft corpus, known as the Minecraft Structured Dialogue Corpus (MSDC), using the discourse theory and annotation principles of SDRT Asher, (1993); Asher and Lascarides, (2003) extended to a multimodal environment, in which both nonlinguistic actions and discourse moves can enter into semantic relations like Elaboration, Correction, and Narration Hunter et al., (2018); Asher et al., (2020). They followed annotation practices given for the STAC corpus Asher et al., (2016).  Thompson et al., (2024) also adapted the parser from  Bennis et al., (2023) to predict discourse structures for the Minecraft corpus with relatively high reliability. Statistics on the MSDC are in Table 1.
@@
Table 1: MDC and MSDC characteristics.
LLMs in robotics Parallel to this work, there has been an increasing amount of research in aiding virtual or real robots with tasks by using LLMs to provide translations from natural language instruction to code that programs the robot to perform the relevant actions Liang et al., (2023); Singh et al., (2023); Yu et al., (2023). This research is directly relevant to our work, as we use LLMs to go from natural language to a pseudo-code of pick and place statements. However, whereas Liang et al., (2023); Singh et al., (2023); Yu et al., (2023) focus on optimizing the translation from instructions, typically one instruction, to various different coding paradigms, we focus on how linguistic and nonlinguistic interactions affect the resulting action sequence. As our results and previous results on the MDC show, producing actions from interactive conversation with frequently underspecified instructions, which are also dependent upon the discourse and nonlinguistic contexts for proper interpretation, is a much more challenging task than translating well crafted unambiguous instructions into code. In addition, we show that to predict a relevant action from the instruction $i_{n+1}$ in the MDC environment, it is not sufficient to use a context with just the penultimate instruction $i_{n}$ and previous action sequence $a_{n}$.
Figure 2: An excerpt from Minecraft Corpus. The Builder interrupts the action sequence by asking a question.
@@ 3 NeBuLa: an LLM for Predicting Action Sequences
We’ve seen that Jayannavar et al., (2020)’s evaluation method for neural agents gives rather poor results. Observations of the results of neural Builder anecdotally yielded no ending configurations that matched those in the gold. The training scheme of Jayannavar et al., (2020) assumes, in effect, that Architect instructions and Builder actions follow one another with regularity. An unfortunate consequence of this assumption is that actions are individuated by the conversational turns that immediately precede and follow them. Jayannavar et al., (2020) initiate a new action sequence whenever there is a linguistic move of any kind. But that’s not realistic, as bits of text don’t always yield a well-formed or even an underspecified instruction. You might have a clarification question from the Builder in between two action sequences that are in fact carrying out one and the same action as in Figure 2. Builders in the MDC frequently ask questions with respect to the initial instruction about the actions they are simultaneously making; answers to those questions may affect the actions, but it doesn’t mean that there are two distinct series of actions pertaining to two distinct instructions, one before the question and its response and one after. In addition, the Builder sometimes starts to build before the instruction sequence is complete; intuitively, the initial actions form a coherent action sequence with the actions that are subsequent to the further instruction. These observations show that the assumptions of Jayannavar et al., (2020) about how actions are individuated are too simple.
@@
Different conversational moves will change and make more precise the shape and position of the structure intended by the initial instruction. Hunter et al., (2018) note that different conversational moves can help conceptualize actions differently. For example, in many Minecraft sessions, an initial instruction gives the Builder an action type that might be realized in many different ways. Something like build a tower of 5 blocks is an action type for which a concrete realization would have to specify the color, perhaps the nature of the building blocks, and a location. As the conversation evolves and unless the Architect corrects their instruction, the type of action to be performed becomes more and more specified.
@@
A simple baseline alternative to the scheme proposed by Jayannavar et al., (2020) that addresses the difficulties we just mentioned is to see how a model performs with the complete prior conversation and action sequences up to the predicted action. This was not an option for Jayannavar et al., (2020)’s model, but more recent LLMs are capable of doing this.
@@
We used Llama-2-7B, Llama-2-13B and Llama-3-8B models to take as context all the conversation and action sequences up to action sequence $a_{n}$ to predict $a_{n}$. We fine-tuned Llama on the MDC’s Jayannavar et al., (2020) training set. All the models were finetuned for $3$ epochs using QLoRA method Dettmers et al., (2023). Table 8 in the appendix provides details of computing resources and the hyperparameters for finetuning. Table 2 shows the net-action F1 scores on the validation and test set of MDC. All the finetuned LLMs significantly improved scores in comparison with the F1 $0.20$ score of Neural Builder Jayannavar et al., (2020). Llama-3-8B essentially doubled the baseline score of $0.20$. In the rest of the paper, we refer to Llama-3-8B finetuned on MDC as NeBuLa. The finetuned model, NeBuLa, is available here [^1] https://huggingface.co/akshay107/NeBuLa.
@@ 4 Using Discourse Structure to Improve NeBuLa
The ideal way to model the instructional interactions is to have two ongoing, interleaved processes that interact and influence each other. On the one hand, there is the evolving conversational structure that helps conceptualize the nonlinguistic actions; on the other, there is the sequence of actions that also affects continuations of the given conversational context.
Using the discourse parser of Thompson et al., (2024), we made a first approximation of these interleaved processes by determining necessary and sufficient situated, conversational conditions for computing instructions. An analysis of the discourse structure in the MSDC shows a large scale pattern of so-called Narrative arcs. These arcs delimit portions of discourse structure linked by Narration relation. Each portion begins with an instruction $i_{n}$ by the Architect, terminates with an action sequence $a_{m}$, and involves a negotiation between Architect and Builder about the action sequence to be performed. The negotiation may be extremely short, where the narrative portion then contains just $i_{n},a_{m}$. On the other hand, it might be complex negotiation involving a number of EDUs related by relations like Elaboration. It may also involve questions of clarification or confirmation question by the Builder, in which case the instruction evolves through the portion. A narrative arc may also involve actions by the Builder that the Architect will correct with a linguistic move that will then result in a nonlinguistic action that revises or corrects the prior actions of the Builder. The end of the negotiation is the action sequence that finally carries out the instructions to the satisfaction of the Architect.
Figure 3 illustrates a narrative arc starting at Architect turn one with a new instruction that results (in green) in an action sequence in Builder turn two. The Builder then asks a complex, alternative question to confirm that this is the right move. The Architect replies to the question, in effect correcting (in red) the Builder’s previous action, which then results in an action sequence in Builder turn four that corrects the previous builder action.
These arcs are relatively self-contained and are recoverable automatically to a relatively high degree by the parser of Thompson et al., (2024). So, instead of providing the entire conversation history as in Section 3 to the model, we provide the world-state at the beginning of the Narrative arc in terms of net place actions, and the discourse within the Narrative arc up to the present instruction $i_{n}$. We finetune Llama-3-8B on MDC training set using this input. We refer to the resultant model as NeBuLa+N (NeBuLa trained on Narrative arcs).
@@
Table 3: Net-Action F1 scores on Minecraft Validation and Test set for predicting action sequences for LLMs. NeBuLa+N refers to NeBuLa trained on narrative arcs. The next two rows look at those 254 examples in the test set where $i_{n}a_{n}i_{n+1}$ has less content than the relevant Narrative arc. NeBuLa+N/N gives score of NeBuLa+N on these samples when world-state and narrative arc is given as input. Similarly, NeBuLa+N/$i_{n}a_{n}i_{n+1}$ gives score of NeBuLa+N on the same samples when world-state and $i_{n}a_{n}i_{n+1}$ is given as input.
Table 3 gives scores on the validation and test set of MDC for NeBuLa+N(arration). From the table, we can see that the scores are comparable with original NeBuLa. This shows that the Narrative arc is sufficient for action prediction. To study whether the Narrative arc is necessary as well, we evaluated NeBuLa+N on those cases where $i_{n}a_{n}i_{n+1}$ had less content than the Narrative arc. There were 254 such cases in MDC test set. For these samples, we looked at performance of NeBuLa+N when worldstate along with the Narrative arc is given as input. This is denoted as NeBuLa+N/N in Table 3. Similarly, we looked at performance of NeBuLa+N when worldstate along with $i_{n}a_{n}i_{n+1}$ is given as input. This is denoted as NeBuLa+N/$i_{n}a_{n}i_{n+1}$ in Table 3. As we can see, the score for NeBuLa+N/$i_{n}a_{n}i_{n+1}$ is significantly lower ($\sim 10\%$) than NeBuLa+N/N. This shows that Narrative arcs are crucial for the task of action prediction.
5 Problems with the Minecraft Corpus

### v2

@@ Abstract
When engaging in collaborative tasks, humans efficiently exploit the semantic structure of a conversation to optimize verbal and nonverbal interactions. But in recent “language to code” or “language to action” models, this information is lacking. We show how incorporating the prior discourse and nonlinguistic context of a conversation situated in a nonlinguistic environment can improve the “language to action” component of such interactions. We finetune an LLM to predict actions based on prior context; our model, Nebula, doubles the net-action F1 score over the baseline on this task of Jayannavar et al., (2020). We also investigate our model’s ability to construct shapes and understand location descriptions using a synthetic dataset.
@@ 1 Introduction
The first factor is the interaction between linguistic and nonlinguistic contexts. Previous work has shown that at least some context is needed to understand and carry out conversationally given instructions Jayannavar et al., (2020). We improve on that work by first establishing a baseline by using the entire exchange up to an instruction $i$ as a context for an LLM to interpret $i$. Our LLM model, Nebula (Neural builder with Llama), trained on the Minecraft Dialogue Corpus (MDC) Narayan-Chen et al., (2019), achieves net-action F1 scores that is almost double of Jayannavar et al., (2020). Using the Minecraft Structured Dialogue dataset (MSDC) Thompson et al., 2024b , which provides semantic relations between MDC dialogue moves and nonlinguistic actions, we show that particular discursive components of the linguistic and nonlinguistic context are necessary and sufficient for the LLM to understand an instruction to the degree provided by the baseline.
@@
Analysing Nebula’s output reveals two other factors that adversely affect its performance. An instruction in the MSDC has two basic components: a description of a shape in terms of four parameters—numbers of components, colors, arrangement and orientation—and the description of a location where the shape should be placed. Human Architects often use analogies to everyday objects that may be challenging to process; in addition, shape descriptions are often underspecified, meaning that one could perform the instruction correctly in various ways. Location descriptions in the Minecraft world are also quite difficult to process and highly underspecified. For example, put a tower in a corner could be correctly located in any of the four corners of the Minecraft board. We address this problem in two ways: first by further finetuning Nebula on a synthetic dataset to improve its performance in building basic shapes and locating them appropriately; and secondly, and more importantly, by revising the evaluation metric used by Jayannavar et al., (2020) to reflect more realistically the semantics of location expressions. We show that, on our synthetic dataset, Nebula achieves high accuracy as per our intuitive metric in performing basic instructions.
The main contributions of this work are: (i) to show that finetuning Llama-3-8B Dubey et al., (2024) on the entire prior history of the action-prediction task Nebula doubles the net-action F1 score on the Minecraft dataset as compared with the Neural Builder of Jayannavar et al., (2020); (ii) to show that training on Narrative arcs achieves comparable results with baseline Nebula, and that training on narrative arcs performs better than the instruction-action-instruction input template used by Jayannavar et al., (2020); (iii) to highlight the drawbacks of the evaluation metric (i.e., net-action F1), and propose and test a new metric on our synthetic datasets.
@@
After some preliminaries and discussion of prior work (Section 2), we present our model, Nebula, and its baseline performance in Section 3, and then a necessary and sufficient discourse feature to get scores equivalent to the baseline in Section 4. In Section 5, we explain several issues associated with the Minecraft corpus. We try to address these issues in Section 6, where we explain our evaluation metric for underspecified instructions, as well as experiments on our synthetic datasets. We conclude in Section 7.
@@ 2 Related Work
MDC Narayan-Chen et al., (2019) construct a corpus of two person dialogues situated in a simulated Minecraft environment. The dialogues record conversations about collaborative tasks, in which an Architect and a Builder cooperate to build sometimes complex 3-dimensional shapes out of blocks of six different colors. The Architect provides instructions, while the Builder is tasked with translating these instructions into actions. The Builder sometimes asks questions, and the Architect may correct themselves or the Builder, or both, concerning both linguistic and nonlinguistic moves. The corpus accurately reflects the variety and complexity of actual cooperative conversation. Details on the MDC are in Table 1. The corpus has also led to related work on conversational interactive agents Kiseleva et al., (2022); Mohanty et al., (2023); Madge and Poesio, (2024).
@@
Instructions to code: Neural Builder and variants The MDC Jayannavar et al., (2020) incentivizes the development of an algorithm that can predict sequences of actions from instructions. The actions involved basic moves of placing or removing blocks from certain positions in the environment. Jayannavar et al., (2020) train a model consisting of a GRU Cho et al., (2014) to handle textual input coupled with a CNN to integrate information from the current state and a GRU to predicted an action sequence. Although they experiment with several training regimes, the best performance comes from one in which a sequence of conversational moves after some action sequence, assumed to be instructions are given to the model, are followed by the next action sequence of the Builder, followed by the next sequence of linguistic moves are input to the model to predict the subsequent action sequence (See Figure 1).
@@
Figure 1: The Neural Builder Jayannavar et al., (2020) takes as input the sequence $i\_n$ $a\_n$ $i\_n+1$ and the worldstate to predict the subsequent action sequence.
@@
The net-action F1 metric evaluates a model’s prediction based on the exact color and coordinate match between the model’s predicted sequence and Builder’s gold action sequence. In general, Jayannavar et al., (2020) show that the problem of predicting action sequences from natural language instructions in naturally occurring dialogue remains extremely challenging. Their Neural Builder has net action F1 of $0.20$ on the MDC test set.
@@
Shi et al., (2022) propose a somewhat different task from Jayannavar et al., (2020); they try to predict when the Builder should execute an action and when they should instead ask for a clarification question. To this end, they annotate all Builder dialogue moves with a taxonomy of dialogue acts. They then specify a single specific action under the execution label instead of a sequence of actions. Thus, their set-up is not directly comparable to that of Jayannavar et al., (2020).
@@
Bonial et al., (2020, 2021) add dialogue acts to Minecraft utterances, but they do not evaluate the effect of these dialogue acts on the Neural Builder’s predictions of actions. Dialogue acts are a partial step towards a full discourse structure: they provide labels for various dialogue moves, but the full discourse structure that we propose to use involves relations between moves. These relations are important as they tell us how to link different parts of, for instance, an instruction into a coherent whole. As we aim to demonstrate in this paper, discourse structure can help to clean up datasets for training and thereby improve training.
@@
MSDC Thompson et al., 2024b provide full discourse annotations for the MDC, known as the Minecraft Structured Dialogue Corpus (MSDC), using the discourse theory and annotation principles of SDRT Asher, (1993); Asher and Lascarides, (2003) extended to a multimodal environment, in which both nonlinguistic actions and discourse moves can enter into semantic relations like Elaboration, Correction, and Narration Hunter et al., (2018); Asher et al., (2020). They follow annotation practices given for the STAC corpus Asher et al., (2016).  Thompson et al., 2024b also adapt the parser from  Bennis et al., (2023) to predict discourse structures for the MDC with relatively high reliability. Statistics on the MSDC are in Table 1.
@@
Table 1: MDC and MSDC characteristics. EDU and EEU refer to elementary discourse unit, and elementary event units respectively.
LLMs in robotics Parallel to this work, there has been an increasing amount of research in aiding virtual or real robots with tasks by using LLMs to provide translations from natural language instruction to code that programs the robot to perform the relevant actions Liang et al., (2023); Singh et al., (2023); Yu et al., (2023). This research is directly relevant to our work, as we use LLMs to go from natural language to a pseudo-code of pick and place statements. However, whereas Liang et al., (2023); Singh et al., (2023); Yu et al., (2023) focus on optimizing the translation from instructions, typically one instruction, to various different coding paradigms, we focus on how linguistic and nonlinguistic interactions affect the resulting action sequence. As our results and previous results on the MDC show, producing actions from interactive conversation with frequently underspecified instructions, which are also dependent upon the discourse and nonlinguistic contexts for proper interpretation, is a much more challenging task than translating well crafted unambiguous instructions into code. In addition, we show that to predict a relevant action from the current instruction $i_{n+1}$ in the MDC environment, it is not sufficient to use a context of just the penultimate instruction $i_{n}$ and previous action sequence $a_{n}$.
Figure 2: An excerpt from MDC. The Builder interrupts the action sequence by asking a question.
@@ 3 Nebula: an LLM for Predicting Action Sequences
We’ve seen that Jayannavar et al., (2020)’s Neural Builder performs poorly as per their evaluation method, i.e., net-action F1. The training scheme of Jayannavar et al., (2020) assumes, in effect, that Architect instructions and the Builder actions which fulfill these instructions follow one another in a regular succession. A consequence of this assumption is that actions are individuated by the conversational turns that immediately precede and follow them, and likewise that a new action sequence is initiated whenever there is a linguistic move of any kind. Neural Builder predicts actions from a preceding context containing an instruction, and the instruction just before it, and a representation of the builder moves in between: $i_{n}a_{n}i_{n+1}$ predicts $a_{n+1}$. Additionally, Neural Builder is provided a worldstate representation (See Figure 1).
@@
But this is not realistic, as these bits of text don’t always yield a well-formed instruction or even an underspecified one. In addition, often actions and instructions occur simultaneously. We might have a clarification question from the Builder in between two action sequences that are in fact carrying out the same action as in Figure 2. Builders in the MDC frequently ask questions with respect to the initial instruction about the actions they are currently carrying out; answers to those questions may affect the actions, but it doesn’t mean that there are two distinct series of actions pertaining to two distinct instructions, one before the question and its response and one after. In addition, the Builder sometimes starts to build before the instruction sequence is complete; intuitively, the initial actions form a coherent action sequence with the actions that are subsequent to the further instruction. These observations show that the assumptions of Jayannavar et al., (2020) about how actions are individuated are too simple.
@@
The shape and position of the structure, intended by the initial instruction, can change or be made more precise by different conversational moves. Hunter et al., (2018) note that different conversational moves can help conceptualize actions differently. For example, in many Minecraft sessions, an initial instruction gives the Builder an action type that might be realized in many different ways. Something like build a tower of 5 blocks is an action type for which a concrete realization would have to specify the color, and perhaps the nature of the building blocks, as well as a location. As the conversation evolves and unless the Architect corrects their instruction, the type of action to be performed becomes more and more specified.
@@
A simple baseline alternative to the scheme proposed by Jayannavar et al., (2020) that addresses these difficulties is to see how a model performs with the complete prior conversation and action sequences up to the predicted action. This was not an option for Jayannavar et al., (2020)’s model, but more recent LLMs are capable of doing this.
@@
We use Llama-2-7B, Llama-2-13B and Llama-3-8B models to take as context all the conversation and action sequences up to action sequence $a_{n}$ to predict $a_{n}$. We finetune Llama on the MDC’s Jayannavar et al., (2020) training set. All the models are finetuned for $3$ epochs using QLoRA method Dettmers et al., (2023). Table 2 shows the net-action F1 scores on the validation and test set of MDC. All the finetuned LLMs significantly improve scores in comparison with the $0.20$ F1 score of Neural Builder Jayannavar et al., (2020). Llama-3-8B essentially doubles the baseline score of $0.20$. In the rest of this paper, we refer to Llama-3-8B finetuned on MDC as Nebula. The finetuned model, Nebula, and the synthetic datasets used in Section 6 are available here [^1] https://huggingface.co/linagora/Nebula. Table 6 in the Appendix provides details of computing resources and the hyperparameters for finetuning.
@@ 4 Using Discourse Structure to Improve Nebula
The ideal way to model collaborative, instructional interactions like those featured in the MDC is to have two simultaneous, interleaved processes that interact with each other. On one hand, there is the evolving conversational structure that conceptualizes the nonlinguistic actions; on the other, there is the sequence of actions that also affects continuations of the conversational structure.
This interleaved process is modeled by the discourse structure provided by the MSDC of Thompson et al., 2024b . The MSDC shows a large scale pattern of Narrative arcs. These arcs delimit portions of discourse structure linked by a Narration relation. Each portion begins with an instruction $i_{n}$ from the Architect, terminates with an action sequence $a_{m}$, and involves a negotiation between the Architect and the Builder about the action sequence to be performed.
The negotiation may be extremely short, where the arc just contains a single instruction and resulting (correct) action $i_{n},a_{m}$. But it can also contain a complex negotiation involving a number of discourse moves that serve to elaborate or clarify the initial instruction. In this way, the content of a single instruction $i_{n}$ may evolve over many discourse moves. Further, if the initial builder actions are incorrect, this will prompt corrective moves by the Architect and subsequent revisions by the Builder, leading to longer collaborative negotiations until the Builder carries out the action sequence that satisfies the Architect’s initial instruction.
@@
Figure 3 illustrates a long negotiation enclosed by a Narrative arc. The instruction in the first turn results (in green) in a Builder action sequence in turn two. The Builder then asks a question to confirm that the actions in turn two were correct. The Architect replies to the question by correcting (in red) the actions in turn two, which then results in a corrective action sequence in turn four.
Guided by the assumption that these Narrative arcs contain all the contextual information—the initial instruction and all subsequent discourse moves—that is needed to predict the correct actions, we use the Narrative arcs provided by the MSDC to assign the previous context to each action sequence in the MDC. Since the arcs are automatically recoverable to a relatively high degree by the parsers discussed in Thompson et al., 2024a ; Thompson et al., 2024b , we make the expedient choice to use the gold arcs to test our assumption.
We finetune Llama-3-8B on the MDC training set using instruction-action pairs, where the instruction is only the conversation within the Narrative arc up to the present instruction $i_{n+1}$ in the pair, not the entire conversation history, which we fed to baseline Nebula. Since Neural Builder supplements attenuated contexts with a worldstate representation (see Section 3) we also include a “worldstate” at the beginning of the Narrative arc in terms of net place actions. We refer to the resulting model as Nebula+N (Nebula trained on Narrative arcs).
@@
Table 3: Net-Action F1 scores on Minecraft validation and test sets for predicting action sequences for LLMs. Nebula+N refers to Nebula trained on Narrative arcs. The next two rows look at those $254$ examples in the test set where $i_{n}a_{n}i_{n+1}$ has less content than the associated Narrative arc. Nebula+N/N gives score of Nebula+N on these samples when worldstate and narrative arc is given as input. Similarly, Nebula+N/$i_{n}a_{n}i_{n+1}$ gives score of Nebula+N on the same samples when worldstate and $i_{n}a_{n}i_{n+1}$ is given as input.
Table 3 shows scores on the validation and test sets of the MDC for Nebula+N. We can see that the scores are comparable with Nebula, $0.38$ compared to $0.39$ F1, (see Llama-3-8B in Table 2). This provides evidence that the discourse information present in a Narrative arc is sufficient for action prediction within that arc.
The majority of the Narrative arcs are shorter than or of equivalent length to their $i_{n}a_{n}i_{n+1}$ counterparts used to train the Neural Builder ($1321$ out of $1575$ samples in the test set). Going on length alone, it is reasonable to assume Narrative arcs and $i_{n}a_{n}i_{n+1}$ are interchangeable. However, if Narrative arcs are not only sufficient but also necessary for action prediction, then in the $254$ samples where Narrative arc is longer than $i_{n}a_{n}i_{n+1}$ we would expect Nebula+N to perform better on those samples when given the Narrative arc. And this is the case.
In Table 3, we compare the performance of Nebula+N when worldstate along with the Narrative arc is given as input (denoted as Nebula+N/N), with Nebula+N when worldstate along with $i_{n}a_{n}i_{n+1}$ is given as input (denoted as Nebula+N/$i_{n}a_{n}i_{n+1}$). As we can see, the score for Nebula+N/$i_{n}a_{n}i_{n+1}$ is considerably lower than Nebula+N/N. This provides evidence that Narrative arcs play a crucial role, in that the information they provide is both necessary and sufficient for action prediction in the MDC.
5 Problems with the MDC